OpenStudy (anonymous):

Divide 36x3 - 96x2 + 48x by 12x. 3x2 - 8x + 4 3x2 + 8x - 4 3x3 - 8x + 4 3x3 + 8x - 4

OpenStudy (mathmate):

Divide each term of the polynomial by 12x, the standard method to divide a polynomial by a monomial. Example, for the first term: \(36x^3 \div 12x =(36/12)\times(x^3\div x)=3x^2 \)

OpenStudy (anonymous):

thanks :) i got it

OpenStudy (mathmate):

You're welcome! :)